# Branching Brownian motion versus Random Energy Model in the supercritical phase: overlap distribution and temperature susceptibility

### v2

@@ 1.1 Motivations
Let us also mention that another important and famous example of log-correlated Gaussian fields is the two-dimensional discrete Gaussian Free Field (GFF),which possesses a complicated (non-hierachical) structure of extreme values, but it turns out to be possible to compare it with that of the branching random walk. By comparison to analoguous results for branching random walks, many deep results have been recently established by Bolthausen, Deuschel and Giacomin [9], Daviaud [23], Arguin and Zindy [4], Bolthausen, Deuschel and Zeitouni [10], Bramson and Zeitouni [16], Bramson, Ding and Zeitouni [15], Biskup and Louidor [6, 7]. We refer to the excellent notes by Biskup [5] for more details.
While from a probabilistic point of view, the difference between BBM and REM is perfectly known, it is not the same when quantities from statistical physics are considered. Indeed it is not clear when the decorations of the log-correlated models are felt at the level of the Gibbs measure. For example, Bonnefont [12] recently proved for the BBM that the distribution of the overlap between two points sampled independently according to Gibbs measures at different temperatures is different than the REM’s one, while it is the same for both models if the two points are sampled at the same temperature. This raises questions about the influence of these decorations. Motivated by the recent work of Derrida and Mottishaw [26], one seeks to compare both models by studying carefuly the overlap distribution in the neighborhood of the critical temperature. Another quantity of interest is the notion of temperature susceptibility introduced and studied for the REM (with exponentially distributed energies) by Sales and Bouchaud [46].
@@ 1.2 Definitions and some results
$\sum_{x\in L_{t}}\delta_{h_{t}(x)-t+\frac{3}{2}\log t-\log(c_{d}D_{\infty})}\xrightarrow[t\to\infty]{{\rm(d)}}\sum_{i,j}\delta_{\xi_{i}+d_{ik}},$ (1.1)
for some positive constant $c_{d}>0$ and where $D_{\infty}\coloneqq\lim_{t\to+\infty}\sum_{x\in L_{t}}(t-h_{t}(x))\,\mathrm{e}^{h_{t}(x)-t}>0$ (almost-surely) is the limiting derivative martingale, $\sum_{i\geq 1}\delta_{\xi_{i}}$ is a Poisson point process on $\mathbb{R}$ with intensity measure $\mathrm{e}^{-x}\mathop{}\mathopen{}\mathrm{d}x$ independent of $(\sum_{k\geq 0}\delta_{d_{ik}})_{i\geq 1}$, which are are i.i.d. copies of a point process $\mathcal{D}$ on $(-\infty,0]$ which has a.s. an atom at 0. The point process $\mathcal{D}$ is called decoration process [^2] We will also consider the case of general decoration processes, see Section 2.2 and use the same notation $(\sum_{k\geq 0}\delta_{d_{ik}})_{i\geq 1}$. If the case of general decoration processes is studied, this will be specified, otherwise it is only the BBM’s decoration process..
@@
In statistical physics, it is common to consider first the partition function $Z_{t,d}(\beta)$ of the model, here the BBM ($\beta$ stands for the inverse-temperature):
$Z_{t,d}(\beta)\coloneqq\sum_{x\in L_{t}}\mathrm{e}^{\beta h_{t}(x)},\qquad\forall\,\beta>0,$
@@
$f_{t,d}(\beta)\coloneqq\frac{1}{t}\,\log Z_{t,d}(\beta),\qquad\forall\,\beta>0.$
@@
It is well known, see [27], that the BBM exhibits a phase transition at the level of the free energy and that this latter is the same as for the REM, namely
@@
$f_{d}(\beta)\coloneqq\lim_{t\to\infty}f_{t,d}(\beta)=\begin{cases}\frac{1+\beta^{2}}{2},&{\rm if}\ \beta<\beta_{c}\coloneqq 1,\\ \beta,&{\rm if}\ \beta\geq\beta_{c},\end{cases}\qquad\text{ a.s. and in $L^{1}$.}$
@@
In particular, the model undergoes freezing above $\beta_{c}$ in the sense that the quantity $f(\beta)/\beta$ is constant. More importantly, one considers the Gibbs measure; this is the random probability measure ${\cal G}_{\beta,t,d}$ on $L_{t}$ defined, at inverse temperature $\beta>0$, by
${\cal G}_{\beta,t,d}(x)\coloneqq\frac{\mathrm{e}^{\beta h_{t}(x)}}{Z_{t,d}(\beta)},\qquad\forall\,x\in L_{t}.$
@@
By design, the Gibbs measure concentrates at low temperature, i.e. when $\beta>\beta_{c}$, on the high points of the Gaussian field. With spin glasses in mind, one also considers the normalized covariance or overlap
@@
$q_{t,d}(x,y)\coloneqq\frac{1}{t}\,\mathbb{E}\mathopen{}\mathclose{{}\left[h_{t}(x)h_{t}(y)}\right]=\frac{1}{t}\sup\mathopen{}\mathclose{{}\left\{s\geq 0\,:\,x,y\text{ share a common ancestor in }L_{s}}\right\},\quad\forall\,x,y\in L_{t}.$
@@
A fundamental object, that records the correlations of high points, is the distribution function of the overlap sampled from the Gibbs measure, i.e. $\mathcal{G}_{\beta,t,d}\otimes\mathcal{G}_{\beta^{\prime},t,d}(q_{t,d}(u,v)\geq a)$ for any $a\in(0,1)$, where $u$ (respectively $v$) is sampled according to $\mathcal{G}_{\beta,t,d}$ (respectively $\mathcal{G}_{\beta^{\prime},t,d}$). Bonnefont [12] recently proved that, if $\beta\leq\beta_{c}$ or $\beta^{\prime}\leq\beta_{c}$, then $\mathcal{G}_{\beta,t,d}\otimes\mathcal{G}_{\beta^{\prime},t,d}(q_{t,d}(u,v)\geq a)$ tends to 00 in $L^{1}$ for all $a\in(0,1)$, while if $\beta>\beta_{c}$ and $\beta^{\prime}>\beta_{c}$, then, for all $a\in(0,1)$,
@@
$\mathcal{G}_{\beta,t,d}\otimes\mathcal{G}_{\beta^{\prime},t,d}(q_{t,d}(u,v)\geq a)\xrightarrow[t\to\infty]{{\rm(d)}}\frac{\sum_{i}\mathopen{}\mathclose{{}\left(\mathrm{e}^{\beta\xi_{i}}\sum_{k}\mathrm{e}^{\beta d_{ik}}}\right)\mathopen{}\mathclose{{}\left(\mathrm{e}^{\beta^{\prime}\xi_{i}}{\sum_{k}\mathrm{e}^{\beta^{\prime}d_{ik}}}}\right)}{\mathopen{}\mathclose{{}\left(\sum_{i,k}\mathrm{e}^{\beta\xi_{i}}\mathrm{e}^{\beta d_{ik}}}\right)\mathopen{}\mathclose{{}\left(\sum_{i,k}\mathrm{e}^{\beta^{\prime}\xi_{i}}\mathrm{e}^{\beta^{\prime}d_{ik}}}\right)}\eqqcolon Q_{d}(\beta,\beta^{\prime}),$ (1.3)
@@
where the $(\xi_{i})_{i}$ and the $(d_{ik})_{i,k}$ were introduced for the description of the limiting extremal process of the BBM, see Equation (1.1). In other words, this result proves the convergence of the pushforward of the measure $\mathcal{G}_{\beta,t,d}\otimes\mathcal{G}_{\beta^{\prime},t,d}$ on $L_{t}^{2}$ by the function $q_{t,d}$, which is a random measure on $[0,1]$. The limit is either $\delta_{0}$ if $\beta\wedge\beta^{\prime}\leq\beta_{c}$, or $(1-Q_{d}(\beta,\beta^{\prime}))\delta_{0}+Q_{d}(\beta,\beta^{\prime})\delta_{1}$ otherwise. Note that, when $\beta>\beta_{c}$ and $\beta^{\prime}>\beta_{c}$, the random variables $(\mathrm{e}^{\beta\xi_{i}}\sum_{k}\mathrm{e}^{\beta d_{ik}}/\sum_{j}\mathrm{e}^{\beta\xi_{j}}\sum_{k}\mathrm{e}^{\beta d_{jk}})_{i\geq 1}$ are the asymptotic weights of the clusters under $\mathcal{G}_{\beta,t,d},$ such that $Q_{d}(\beta,\beta^{\prime})$ is simply the probability of choosing two points in the same cluster (when they are chosen proportionally to their Gibbs weights with inverse temperature $\beta$ and $\beta^{\prime}$ respectively).
@@
It is well known that in the case $\beta^{\prime}=\beta$, the random variables $Q_{d}(\beta,\beta)$ and $Q(\beta,\beta)$ have the same distribution (see [43, Equation (1.2)] for more details). Therefore a natural question one may ask is whether $Q_{d}(\beta,\beta^{\prime})$ and $Q(\beta,\beta^{\prime})$ still have the same distribution when $\beta\neq\beta^{\prime}>\beta_{c}$. Following the work by Pain and Zindy [43] for the two-dimensional discrete Gaussian free field, Bonnefont [12] proved, for the BBM, that
$F_{d}(\beta,\beta^{\prime})\coloneqq\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{d}(\beta,\beta^{\prime})}\right]<\mathbb{E}\mathopen{}\mathclose{{}\left[Q(\beta,\beta^{\prime})}\right]\eqqcolon F(\beta,\beta^{\prime}),$ (1.4)
meaning that the answer is negative. In this paper, our aim is to study and compare the functions $\beta\mapsto F_{d}(\beta,\beta^{\prime})$ and $\beta\mapsto F(\beta,\beta^{\prime})$, when $\beta^{\prime}>\beta_{c}$ is fixed. We will specially focus on their behavior when $\beta$ tends to $\beta_{c}^{+}$.
@@
$Z(\beta)\coloneqq\sum_{i\geq 1}\mathrm{e}^{\beta\xi_{i}},\qquad Z_{d}(\beta)\coloneqq\sum_{i\geq 1}\sum_{k\geq 0}\mathrm{e}^{\beta(\xi_{i}+d_{ik})}.$ (1.5)
@@
$\mathbf{C}(\beta,\beta^{\prime})\coloneqq\frac{\operatorname{Cov}(\log Z(\beta),\log Z(\beta^{\prime}))}{\sigma(\log Z(\beta))\,\sigma(\log Z(\beta^{\prime}))},\qquad\mathbf{C}_{d}(\beta,\beta^{\prime})\coloneqq\frac{\operatorname{Cov}(\log Z_{d}(\beta),\log Z_{d}(\beta^{\prime}))}{\sigma(\log Z_{d}(\beta))\,\sigma(\log Z_{d}(\beta^{\prime}))},$
where $\sigma(X)\coloneqq\sqrt{\operatorname{Var}(X)}$. The susceptibility to temperature we are interested in is defined by Sales and Bouchaud [46] as the coefficient $\kappa_{d}(\beta)$ for the BBM and $\kappa(\beta)$ for the REM such that
@@
$\mathbf{C}_{d}(\beta,\beta+h)$ $=1-\kappa_{d}(\beta)\,h^{2}+o(h^{2}),\qquad h\rightarrow 0.$
@@ 1.3 Main results
Our first results concern the behavior of the mean overlap $F(\beta,\beta^{\prime})$ or $F_{d}(\beta,\beta^{\prime})$ at two supercritical inverse temperatures $\beta$ and $\beta^{\prime}$. Recall from Equation (1.4) that it has already been proved that $F_{d}(\beta,\beta^{\prime})<F(\beta,\beta^{\prime})$. Our goal here is to quantify this difference in the regime where $\beta^{\prime}$ is fixed but $\beta$ approaches the critical inverse temperature $\beta_{c}=1$.
We start with the REM case. It is not hard to see that the function $\beta\mapsto F(\beta,\beta^{\prime})$ has an infinite right-derivative at 1, see Remark 3.10. We establish a sharper estimate in the following result.
@@
$F(\beta,\beta^{\prime})=(\beta-1)\log\frac{1}{\beta-1}+O(\beta-1).$
We compare this behavior to the one arising for the BBM. Surprisingly, the decorations have the effect of smoothing the function $\beta\mapsto F_{d}(\beta,\beta^{\prime})$ at criticality: more precisely, $F_{d}(\beta,\beta^{\prime})$ is of order $(\beta-1)$ as $\beta\downarrow 1$, as shown in the following result. It would be interesting to obtain an asymptotic equivalent instead, which would prove that $\beta\mapsto F_{d}(\beta,\beta^{\prime})$ is right-differentiable at $\beta=1$, but this seems out of reach with the techniques used here.
@@
$0<\liminf_{\beta\downarrow 1}\,\frac{F_{d}(\beta,\beta^{\prime})}{\beta-1}\leq\limsup_{\beta\downarrow 1}\,\frac{F_{d}(\beta,\beta^{\prime})}{\beta-1}<+\infty.$
@@
  • If the sum $\sum_{k\geq 0}\mathrm{e}^{\beta d_{k}}$ has bounded $(1+\varepsilon)$-moment as $\beta\downarrow 1$, then the behavior of $F_{d}(\beta,\beta^{\prime})$ is the same as for the REM in Theorem 1.1, see Corollary 3.2.
  • If the $\gamma$-th moment of the sum $\sum_{k\geq 0}\mathrm{e}^{\beta d_{k}}$ explodes like $(\beta-1)^{-\psi(\gamma)}$ as $\beta\downarrow 1$, then different behaviors appear depending on the value of $\psi^{\prime}(1)-\psi(1)$: the larger this value—in other words, the more moments of $\sum_{k\geq 0}\mathrm{e}^{\beta d_{k}}$ are governed by rare events—the faster $F_{d}(\beta,\beta^{\prime})$ decreases as $\beta\downarrow 1$. More precisely, when $\psi^{\prime}(1)-\psi(1)<1$, $F_{d}(\beta,\beta^{\prime})$ is still of order $(\beta-1)\log\frac{1}{\beta-1}$ like in the REM case, but possibly with another multiplicative constant. At the critical value $\psi^{\prime}(1)-\psi(1)=1$, the first order constant vanishes and $F_{d}(\beta,\beta^{\prime})$ becomes of order $(\beta-1)$. Above the critical value, $F_{d}(\beta,\beta^{\prime})$ decreases faster than $(\beta-1)^{1+\eta}$ for $\eta>0$ in some explicit interval, see Corollary 3.5 for a precise statement.

### v3

@@ 1.1 Motivations
Let us also mention that another important and famous example of log-correlated Gaussian fields is the two-dimensional discrete Gaussian Free Field (GFF),which possesses a complicated (non-hierachical) structure of extreme values, but it turns out to be possible to compare it with that of the branching random walk. By comparison to analoguous results for branching random walks, many deep results have been recently established by Bolthausen, Deuschel and Giacomin [11], Daviaud [28], Arguin and Zindy [4, 5], Bolthausen, Deuschel and Zeitouni [12], Bramson and Zeitouni [21], Bramson, Ding and Zeitouni [20], Biskup and Louidor [8, 9]. We refer to the excellent notes by Biskup [7] for more details.
While from a probabilistic point of view, the difference between BBM and REM is perfectly known, it is not the same when quantities from statistical physics are considered. Indeed it is not clear when the decorations of the log-correlated models are felt at the level of the Gibbs measure. For example, motivated by the question of temperature chaos for spin glasses (see the introduction of [48] for more comments on this), Bonnefont [14] recently proved for the BBM that the distribution of the overlap between two points sampled independently according to Gibbs measures at different temperatures is different than the REM’s one, while it is the same for both models if the two points are sampled at the same temperature. This raises questions about the influence of these decorations. Motivated by the recent work of Derrida and Mottishaw [31], we seek to compare both models by studying carefuly the overlap distribution in the neighborhood of the critical temperature. Another quantity of interest is the notion of temperature susceptibility introduced and studied for the REM (with exponentially distributed energies) by Sales and Bouchaud [52].
@@ 1.2 Definitions and some results
$\sum_{x\in L_{t}}\delta_{h_{t}(x)-t+\frac{3}{2}\log t-\log(c_{\mathrm{d}}D_{\infty})}\xrightarrow[t\to\infty]{{\rm(d)}}\sum_{i,j}\delta_{\xi_{i}+d_{ik}},$ (1.1)
for some constant $c_{\mathrm{d}}>0$ and where $D_{\infty}\coloneqq\lim_{t\to+\infty}\sum_{x\in L_{t}}(t-h_{t}(x))\,\mathrm{e}^{h_{t}(x)-t}>0$ (almost-surely) is the limiting derivative martingale, $\sum_{i\geq 1}\delta_{\xi_{i}}$ is a Poisson point process on $\mathbb{R}$ with intensity measure $\mathrm{e}^{-x}\mathop{}\mathopen{}\mathrm{d}x$ independent of $(\sum_{k\geq 0}\delta_{d_{ik}})_{i\geq 1}$, which are are i.i.d. copies of a point process $\mathcal{D}$ on $(-\infty,0]$ which has a.s. an atom at 0. The point process $\mathcal{D}$ is called decoration process [^2] We will also consider the case of general decoration processes, see Section 2.2 and use the same notation $(\sum_{k\geq 0}\delta_{d_{ik}})_{i\geq 1}$ for general decoration processes. For clarity, one shall make precise (between brackets), in every statement/result, when one treats the BBM case or the general decorated case..
@@
In statistical physics, it is common to consider first the partition function $Z_{t,{\mathrm{d}}}(\beta)$ of the model, here the BBM ($\beta$ stands for the inverse-temperature):
$Z_{t,{\mathrm{d}}}(\beta)\coloneqq\sum_{x\in L_{t}}\mathrm{e}^{\beta h_{t}(x)},\qquad\forall\,\beta>0,$
@@
$f_{t,{\mathrm{d}}}(\beta)\coloneqq\frac{1}{t}\,\log Z_{t,{\mathrm{d}}}(\beta),\qquad\forall\,\beta>0.$
@@
The subscript “d” stands for “decorated” and is here to highlight the difference with the REM (for which similar quantities are written without the subscript). It is well known, see [32], that the BBM exhibits a phase transition at the level of the free energy and that this latter is the same as for the REM, namely
@@
$f_{\mathrm{d}}(\beta)\coloneqq\lim_{t\to\infty}f_{t,{\mathrm{d}}}(\beta)=\begin{cases}\frac{1+\beta^{2}}{2},&{\rm if}\ \beta<\beta_{c}\coloneqq 1,\\ \beta,&{\rm if}\ \beta\geq\beta_{c},\end{cases}\qquad\text{ a.s. and in $L^{1}$.}$
@@
In particular, the model undergoes freezing above $\beta_{c}$ in the sense that the quantity $f_{\mathrm{d}}(\beta)/\beta$ is constant. More importantly, one considers the Gibbs measure; this is the random probability measure ${\cal G}_{\beta,t,{\mathrm{d}}}$ on $L_{t}$ defined, at inverse temperature $\beta>0$, by
${\cal G}_{\beta,t,{\mathrm{d}}}(x)\coloneqq\frac{\mathrm{e}^{\beta h_{t}(x)}}{Z_{t,{\mathrm{d}}}(\beta)},\qquad\forall\,x\in L_{t}.$
@@
By design, the Gibbs measure concentrates at low temperature, i.e. when $\beta>\beta_{c}$, on the high points of the Gaussian field. With spin glasses in mind, one also considers the normalized covariance or overlap between two particles $x,y\in L_{t}$ by
@@
$q_{t,{\mathrm{d}}}(x,y)\coloneqq\frac{1}{t}\,\mathbb{E}\mathopen{}\mathclose{{}\left[h_{t}(x)h_{t}(y)}\right]=\frac{1}{t}\sup\mathopen{}\mathclose{{}\left\{s\geq 0\,:\,x,y\text{ share a common ancestor in }L_{s}}\right\}\,.$
@@
A fundamental object, that records the correlations of high points, is the distribution function of the overlap sampled from the Gibbs measure, i.e. $\mathcal{G}_{\beta,t,{\mathrm{d}}}\otimes\mathcal{G}_{\beta^{\prime},t,{\mathrm{d}}}(q_{t,{\mathrm{d}}}(u,v)\geq a)$ for any $a\in(0,1)$, where $u$ (respectively $v$) is sampled according to $\mathcal{G}_{\beta,t,{\mathrm{d}}}$ (respectively $\mathcal{G}_{\beta^{\prime},t,{\mathrm{d}}}$). Bonnefont [14] recently proved that, if $\beta\leq\beta_{c}$ or $\beta^{\prime}\leq\beta_{c}$, then $\mathcal{G}_{\beta,t,{\mathrm{d}}}\otimes\mathcal{G}_{\beta^{\prime},t,{\mathrm{d}}}(q_{t,{\mathrm{d}}}(u,v)\geq a)$ tends to 00 in $L^{1}$ for all $a\in(0,1)$, while if $\beta>\beta_{c}$ and $\beta^{\prime}>\beta_{c}$, then, for all $a\in(0,1)$,
@@
$\mathcal{G}_{\beta,t,{\mathrm{d}}}\otimes\mathcal{G}_{\beta^{\prime},t,{\mathrm{d}}}(q_{t,{\mathrm{d}}}(u,v)\geq a)\xrightarrow[t\to\infty]{{\rm(d)}}\frac{\sum_{i}\mathopen{}\mathclose{{}\left(\mathrm{e}^{\beta\xi_{i}}\sum_{k}\mathrm{e}^{\beta d_{ik}}}\right)\mathopen{}\mathclose{{}\left(\mathrm{e}^{\beta^{\prime}\xi_{i}}{\sum_{k}\mathrm{e}^{\beta^{\prime}d_{ik}}}}\right)}{\mathopen{}\mathclose{{}\left(\sum_{i,k}\mathrm{e}^{\beta\xi_{i}}\mathrm{e}^{\beta d_{ik}}}\right)\mathopen{}\mathclose{{}\left(\sum_{i,k}\mathrm{e}^{\beta^{\prime}\xi_{i}}\mathrm{e}^{\beta^{\prime}d_{ik}}}\right)}\eqqcolon Q_{\mathrm{d}}(\beta,\beta^{\prime}),$ (1.3)
@@
where the $(\xi_{i})_{i}$ and the $(d_{ik})_{i,k}$ were introduced for the description of the limiting extremal process of the BBM, see Equation (1.1). In other words, this result proves the convergence of the pushforward of the measure $\mathcal{G}_{\beta,t,{\mathrm{d}}}\otimes\mathcal{G}_{\beta^{\prime},t,{\mathrm{d}}}$ on $L_{t}^{2}$ by the function $q_{t,{\mathrm{d}}}$, which is a random measure on $[0,1]$. The limit is either $\delta_{0}$ if $\beta\wedge\beta^{\prime}\leq\beta_{c}$, or $(1-Q_{\mathrm{d}}(\beta,\beta^{\prime}))\delta_{0}+Q_{\mathrm{d}}(\beta,\beta^{\prime})\delta_{1}$ otherwise. Note that, when $\beta>\beta_{c}$ and $\beta^{\prime}>\beta_{c}$, the random variables $(\mathrm{e}^{\beta\xi_{i}}\sum_{k}\mathrm{e}^{\beta d_{ik}}/\sum_{j}\mathrm{e}^{\beta\xi_{j}}\sum_{k}\mathrm{e}^{\beta d_{jk}})_{i\geq 1}$ are the asymptotic weights of the clusters under $\mathcal{G}_{\beta,t,{\mathrm{d}}},$ such that $Q_{\mathrm{d}}(\beta,\beta^{\prime})$ is simply the probability of choosing two points in the same cluster (when they are chosen proportionally to their Gibbs weights with inverse temperature $\beta$ and $\beta^{\prime}$ respectively). Observe that, in particular, this result implies absence of temperature chaos for the BBM.
@@
It is well known that in the case $\beta^{\prime}=\beta$, the random variables $Q_{\mathrm{d}}(\beta,\beta)$ and $Q(\beta,\beta)$ have the same distribution (see [48, Equation (1.2)] for more details). Therefore a natural question one may ask is whether $Q_{\mathrm{d}}(\beta,\beta^{\prime})$ and $Q(\beta,\beta^{\prime})$ still have the same distribution when $\beta\neq\beta^{\prime}>\beta_{c}$. Following the work by Pain and Zindy [48] for the two-dimensional discrete Gaussian free field, Bonnefont [14] proved, for the BBM, that
$\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]<\mathbb{E}\mathopen{}\mathclose{{}\left[Q(\beta,\beta^{\prime})}\right],$ (1.4)
meaning that the answer is negative. In this paper, our aim is to study and compare the functions $\beta\mapsto\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]$ and $\beta\mapsto\mathbb{E}\mathopen{}\mathclose{{}\left[Q(\beta,\beta^{\prime})}\right]$, when $\beta^{\prime}>\beta_{c}$ is fixed. We will specially focus on their behavior when $\beta$ tends to $\beta_{c}^{+}$.
@@
$Z(\beta)\coloneqq\sum_{i\geq 1}\mathrm{e}^{\beta\xi_{i}},\qquad Z_{\mathrm{d}}(\beta)\coloneqq\sum_{i\geq 1}\sum_{k\geq 0}\mathrm{e}^{\beta(\xi_{i}+d_{ik})}.$ (1.5)
@@
$\mathbf{C}(\beta,\beta^{\prime})\coloneqq\frac{\operatorname{Cov}(\log Z(\beta),\log Z(\beta^{\prime}))}{\sigma(\log Z(\beta))\,\sigma(\log Z(\beta^{\prime}))},\qquad\mathbf{C}_{\mathrm{d}}(\beta,\beta^{\prime})\coloneqq\frac{\operatorname{Cov}(\log Z_{\mathrm{d}}(\beta),\log Z_{\mathrm{d}}(\beta^{\prime}))}{\sigma(\log Z_{\mathrm{d}}(\beta))\,\sigma(\log Z_{\mathrm{d}}(\beta^{\prime}))},$
where $\sigma(X)\coloneqq\sqrt{\operatorname{Var}(X)}$. The susceptibility to temperature we are interested in is defined by Sales and Bouchaud [52] as the coefficient $\kappa_{\mathrm{d}}(\beta)$ for the BBM and $\kappa(\beta)$ for the REM such that
@@
$\mathbf{C}_{\mathrm{d}}(\beta,\beta+h)$ $=1-\kappa_{\mathrm{d}}(\beta)\,h^{2}+o(h^{2}),\qquad h\rightarrow 0.$
@@
Let us also recall that, for readability, we use the same notation $(\sum_{k\geq 0}\delta_{d_{ik}})_{i\geq 1}$ for the decoration processes of the BBM and of the general decorated case; but, for clarity, it will be precised (between brackets), for every statement/result, when one treats the BBM case or the general decorated case.
@@ 1.3 Main results
Our first results concern the behavior of the mean overlap $\mathbb{E}\mathopen{}\mathclose{{}\left[Q(\beta,\beta^{\prime})}\right]$ or $\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]$ at two supercritical inverse temperatures $\beta$ and $\beta^{\prime}$. Recall from Equation (1.4) that it has already been proved that $\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]<\mathbb{E}\mathopen{}\mathclose{{}\left[Q(\beta,\beta^{\prime})}\right]$. Our goal here is to quantify this difference in the regime where $\beta^{\prime}$ is fixed but $\beta$ approaches the critical inverse temperature $\beta_{c}=1$.
We start with the REM case. It is not hard to see that the function $\beta\mapsto\mathbb{E}\mathopen{}\mathclose{{}\left[Q(\beta,\beta^{\prime})}\right]$ has an infinite right-derivative at 1, see Remark 3.10. We establish a sharper estimate in the following result.
@@
$\mathbb{E}\mathopen{}\mathclose{{}\left[Q(\beta,\beta^{\prime})}\right]=(\beta-1)\log\frac{1}{\beta-1}+O(\beta-1).$
We compare this behavior to the one arising for the BBM. Surprisingly, the decorations have the effect of smoothing the function $\beta\mapsto\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]$ at criticality: more precisely, $\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]$ is of order $(\beta-1)$ as $\beta\downarrow 1$, as shown in the following result. It would be interesting to obtain an asymptotic equivalent instead, which would prove that $\beta\mapsto\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]$ is right-differentiable at $\beta=1$, but this seems out of reach with the techniques used here.
@@
$0<\liminf_{\beta\downarrow 1}\,\frac{\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]}{\beta-1}\leq\limsup_{\beta\downarrow 1}\,\frac{\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]}{\beta-1}<+\infty.$
@@
  • If the sum $\sum_{k\geq 0}\mathrm{e}^{\beta d_{k}}$ has bounded $(1+\varepsilon)$-moment as $\beta\downarrow 1$, then the behavior of $\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]$ is the same as for the REM in Theorem 1.1, see Corollary 3.2.
  • If the $\gamma$-th moment of the sum $\sum_{k\geq 0}\mathrm{e}^{\beta d_{k}}$ explodes like $(\beta-1)^{-\psi(\gamma)}$ as $\beta\downarrow 1$, then different behaviors appear depending on the value of $\psi^{\prime}(1)-\psi(1)$: the larger this value—in other words, the more moments of $\sum_{k\geq 0}\mathrm{e}^{\beta d_{k}}$ are governed by rare events—the faster $\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]$ decreases as $\beta\downarrow 1$. More precisely, when $\psi^{\prime}(1)-\psi(1)<1$, $\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]$ is still of order $(\beta-1)\log\frac{1}{\beta-1}$ like in the REM case, but possibly with another multiplicative constant. At the critical value $\psi^{\prime}(1)-\psi(1)=1$, the first order constant vanishes and $\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]$ becomes of order $(\beta-1)$. Above the critical value, $\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]$ decreases faster than $(\beta-1)^{1+\eta}$ for $\eta>0$ in some explicit interval, see Corollary 3.5 for a precise statement.
@@
We conclude this discussion by two open questions. Firstly, we wonder if the limit in Theorem 1.2 actually exists, or in other words:
Open question 1.3 (BBM case).
For any $\beta^{\prime}>1$, is the function $\beta\in(1,\infty)\mapsto\mathbb{E}[Q_{\mathrm{d}}(\beta,\beta^{\prime})]$ right-differentiable at $1$?
Our results compare two 1-RSB systems, but, as suggested by an anonymous referee, it would be interesting to compare systems with more levels of replica symmetry breaking. More precisely, a $k$-GREM (that is with a tree structure of height $k+1$), where each vertex has $\mathrm{e}^{t/(2k)}$ children and with energy increments $\mathcal{N}(0,\sigma_{\ell}^{2}t)$ at level $\ell$, is a $k$-RSB system if $\sigma_{1}>\dots>\sigma_{k}$, see [16] for details. It can be compared to a $k$-speed BBM, where particles move with variance $\sigma_{\ell}^{2}$ on $[(\ell-1)t,\ell t]$ for a given time horizon $t$, see [15] for a description of its extremal process when $k=2$. With these definitions, the one-temperature overlap distribution is the same for both models in the limit $t\to\infty$, but the two-temperature overlap distribution should be different. It could be compared in an abstract manner in the spirit of [48, 14], or more quantitatively in the near-critical regime as in this paper.
Open question 1.4.
Compare the two-temperature overlap distributions of the $k$-GREM and of the $k$-speed BBM.